Fulton Street Market reported the following for the sales of its only product:

TOTAL PER UNIT
SALES $31,500 $4.50
VARIABLE EXPENSES 9,450 1.35
CONTRIBUTION MARGIN 22,050 $3.15
FIXED EXPENSES 3,000 -
NET OPERATING INCOME 19,050 -

The company would like to increase its selling price by 50 cents per unit, and feels this would decrease unit sales volume by 10%. Should the company increase the price, and what will the effect on net income be>

A. Yes; $3,500 increase

B.Yes; $945 increase

C.No; $945 decrease

D.No; $1,006 decrease

Be sure to show all work!
Answer #1

Selling Price increased by 50% = 4.50+0.50 =5

Sales units Reduced by 10% =(31500/4.50) *90% = 6,300 Units

New Operating income = 6,300*(5.00-1.35)-3000= $19,995

Increase in Operating income = $19,995-19,050 = $945

Answer : B) YES ; $945 Increase